# Scalable learning of potentials to predict time-dependent Hartree-Fock dynamics

### v1

@@ I Introduction
$\imath\frac{d\Psi(\mathbf{r},t)}{dt}=\widehat{H}(\mathbf{r},t)\Psi(\mathbf{r},t).$ (1)
@@
In principle, time-dependent density functional theory (TDDFT) enables one to compute the reduced electron density $\rho$, without first solving the TDSE for $\Psi$ (Ullrich, 2012). However, in TDDFT, the density-dependent exchange-correlation potential (itself a part of the overall inter-electronic potential) is not known and must be approximated (Gross and Maitra, 2012; Maitra, 2016). TDDFT is widely used to simulate electron dynamics and excited state properties, but in some cases the approximations employed for the exchange-correlation potential lead to inaccurate electron dynamics and charge transfer (Elliott et al., 2012; Rappoport and Hutter, 2012; Habenicht et al., 2014; Provorse, Habenicht, and Isborn, 2015; Lacombe and Maitra, 2020; Ranka and Isborn, 2023). Learning a more accurate Hamiltonian, including the unknown density-dependent electron potential of TDDFT, would lead to advances in the simulation of the molecular properties and excited state dynamics.
@@
At the time of writing, there are only three works we know of that seek to learn the TDDFT exchange-correlation potential from data consisting of temporal snapshots of the electron density (Suzuki, Nagai, and Haruyama, 2020; Bhat et al., 2022; Yang and Whitfield, 2023). These works focus on one-dimensional model problems; extending the methods to three-dimensional systems of chemical/physical interest will be a large-scale undertaking from both a theoretical and computational point of view. We lack an understanding of questions such as: (i) how should we form our training sets? (ii) how should we construct our model potentials? and (iii) how should we quantify errors? The literature on learning potentials for quantum molecular systems is tilted heavily toward static, time-independent problems (Snyder et al., 2012; Bartók et al., 2017; Chandrasekaran et al., 2019; Schleder et al., 2019; Jørgensen et al., 2019; Smith et al., 2019; Ceriotti, 2019) and offers little guidance on the answers to these questions, especially for molecular systems with, e.g., dozens of electrons.
@@
The broad goal of this paper is to extend our understanding of how to learn potentials from time series observations of quantum systems. We develop a learning framework for dynamics governed by the time-dependent Hartree-Fock (TDHF) equation. We focus on TDHF for four reasons. First, for TDHF, we know a priori how to compute the entire Hamiltonian, including the potential term we seek to learn. Hence we can compare our learned potentials against the ground truth. Second, TDHF is numerically tractable even for large molecular systems, so there is no a priori obstacle that stops us from generating training data. Third, Hartree-Fock theory incorporates an exact treatment of electron exchange. Consequently, the inter-electronic potential is purely a function of a $2$-electron integral tensor $(ij|kl)$—see (5)—that enjoys an $8$-fold permutation symmetry. Though existing density functionals that approximate electron exchange based on the local density do not have this symmetry, we expect an exact exchange-correlation functional to have symmetries similar to what we see in Hartree-Fock theory. This motivates studying how to incorporate symmetries into a machine learning model of a potential. Fourth, Hartree-Fock is itself the first model (approximating $\Psi$ using a single Slater determinant) in a sequence of increasingly accurate models that culminates in full configuration interaction (in which $\Psi$ is approximated using a linear combination of all possible Slater determinants). Time-dependent full configuration interaction yields numerical solutions of (1) that are exact within a finite-dimensional basis set (McWeeny, 1989).
@@
Therefore, by studying and solving a machine learning problem for TDHF, we aim to produce insights that can be abstracted and generalized to learning problems that involve more accurate models for quantum dynamics. The present study aims to guide future efforts to learn a priori unknown terms that play important roles in time-dependent equations of motion. This includes, but is not limited to, learning the exchange-correlation potential for TDDFT (Suzuki, Nagai, and Haruyama, 2020; Bhat et al., 2022; Yang and Whitfield, 2023), closure terms for time-dependent reduced density matrix functional theory and similar theories (Blum, 2012; Akbari et al., 2012; Ferré, Filatov, and Huix-Rotllant, 2016; Bhat et al., 2024), and surrogate models for the Kadanoff-Baym equations (Bassi et al., 2023).
@@
can be derived from the TDSE by approximating $\Psi$ as a single Slater determinant computed from a finite set of spin orbitals (Szabo and Ostlund, 2012). Here $P$ is the $1$-electron reduced density matrix corresponding to $\Psi$, while the Hamiltonian matrix $H$ is the operator $\widehat{H}$ from (1) expressed in a finite-dimensional basis (McLachlan and Ball, 1964; McWeeny, 1989). Note that (2) only holds if one expresses $P$ and $H$ in orthonormal bases. The full, time-dependent Hamiltonian in (2) is
@@
On the right-hand side, the field-free Hamiltonian $H(P(t))$ includes the kinetic energy of the electrons, the potential due to interaction with the fixed nuclei, and the density dependent inter-electronic repulsion term, also known as the Fock matrix in Hartree-Fock theory. In (3), $V_{\text{ext}}(t)$ is an optional, external, time-dependent potential, e.g., an applied electric field. In our prior work (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022), we initiated the project of learning a model $\widetilde{H}(P)$ of the true field-free Hamiltonian $H(P)$ from time series observations $\{P(t_{j})\}$ of electron density matrices.
@@
In the present work, we move beyond prior work in three ways. First, we pay close attention to scalability to systems such as paranitroaniline ($\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$) in the STO-3G basis set, where the density matrices $P(t)$ are of size $60\times 60$. For $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$, the methods developed in this paper enable learning an inter-electronic potential parameterized by a tensor with more than $1.6$ million entries, with which we can propagate electron dynamics with an overall error on the order of $10^{-7}$—see Figure 4. This is six orders of magnitude better than what we can achieve using methods from our prior work. We are unaware of other methods that can scale to this system size. In the present work, we propose three models and derive expressions for Jacobian-vector products that enable matrix-free, iterative training. We formulate our models in such a way that they preserve symmetries using linear algebra. The combination of these techniques leads to scalability.
Second, in the present work, we explore the generation of training sets that are both highly accurate and highly informative for machine learning purposes. We employ a higher-order numerical method that, for the largest molecular systems we study, avoids spurious oscillations. Using this method, we construct a training set consisting of an ensemble of field-free trajectories. The ensemble explores more of phase space, passing near more minima, maxima, and saddle points that characterize the true Hamiltonian. As we will show, models trained on this ensemble have superior predictive power as measured by the ability to generate field-on trajectories $\widetilde{P}(t)$ that match true field-on trajectories $P(t)$.
Third, at the core of the present work lies the idea of building physics—in the form of symmetries—into our machine learning models. Two of the models we propose preserve Hermitian symmetry of the Hamiltonian $\widetilde{H}(P)$. Another preserves a deeper $8$-fold permutation symmetry that underpins the inter-electronic term in the true field-free Hamiltonian $H(P)$. We algorithmically construct a basis for the space of all $4$-index tensors that satisfy this symmetry. We express our model using this basis, thereby ensuring that the symmetry is preserved. Our methods can be generalized to other tensors, symmetries, and systems.
@@ II.1 Choice of initial condition and time-stepping method
In order to train and test our models, we must have on hand a reliable set of electron dynamics trajectories. We now describe how we generate this data.
For a particular molecule and basis set, we apply standard electronic structure methods to compute the ground truth field-free Hamiltonian $H(P)$ and an initial condition $P(0)$ with an applied delta-kick perturbation. Before doing anything further, we evolve the density matrix forward in time to $P(2\delta t)$ with time step $\delta t=8.268\times 10^{-2}$ a.u. We then relabel this density matrix $P(2\delta t)$ as $P(0)$ and use this as our initial condition for all subsequent modeling/testing. We use a modified version of the Gaussian electronic structure code (Frisch, Trucks, and et. al, 2018) to compute the Hamiltonian $H(P)$ and the matrices $P(0)$ through $P(2\delta t)$.
To understand why we relabel $P(2\delta t)$ as $P(0)$, note that in our past (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022) and present work, our learning procedures depend on first estimating $dP/dt$ from $P(t)$ and then fitting $idP/dt$ to $[H,P]$. That is, we minimize the residual between the left- and right-hand sides of (2). As expected given the nature of the Dirac delta, our numerical estimates of $dP/dt$ are much more accurate if we move past the time of the initial delta-kick perturbation. This improves all learning procedures we have tested.
Equipped with $H(P)$ and $P(0)$, our task is to numerically integrate (2) to produce a trajectory $\{P(t_{j})\}$ of electron density matrices on a temporal grid $t_{j}=j\Delta t$ for $j=0,\ldots,J$. In our prior work (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022), we employed the second-order modified midpoint unitary transformation (MMUT) method (Li et al., 2005a). In the present work, for reasons we now explain, we employ a fourth-order scheme (Casas and Iserles, 2006a) that we abbreviate as CI4. We detail both schemes in Supplementary Information (SI) Section V.1. Note that the CI4 method is based on a fourth-order Magnus expansion. In a comparative study of time-stepping methods for the time-dependent Kohn-Sham equations, a fourth-order Magnus integrator emerged as the clear winner (Gémez Pueyo et al., 2018).
@@
Therefore, for the remainder of this work, for each molecular system (choice of molecule plus basis set), we use the CI4 (Casas and Iserles, 2006a) scheme to solve (2). This includes generation of training and test data using either ground truth Hamiltonians $H(P)$ or learned Hamiltonians $\widetilde{H}(P)$.
@@ II.2 Linear models
We use $\widetilde{H}_{1}(P)$ to denote a model of the density-dependent, inter-electronic potential in (9). We present three scalable $\widetilde{H}_{1}(P)$, in increasing order of mathematical complexity. All three are linear in the entries of $P$. All three preserve Hermitian symmetry. The third model preserves an additional $8$-fold permutation symmetry that we detail below.
@@ II.2.1 Tied Regression
instead of (11b); for this model, the task is to learn both $\bm{\beta}$ and $\bm{\gamma}$. Model (11) effectively ties the parameters $\bm{\beta}$ and $\bm{\gamma}$ to one another (via $\bm{\beta}=\bm{\gamma}$), reducing the number of real parameters from $2N^{4}$ to $N^{4}$, and thus reducing training times. In SI Section V.2.1, we show that there exists $\bm{\beta}$ such that the tied model (11) matches the ground truth Hamiltonian (in the CO basis). Thus there is nothing gained by considering a model in which (12) replaces (11b).
@@ II.2.2 Regression with Hermitian Representation
$\sum_{i,j}P^{R}_{ij}\beta_{ijkl}=[P^{R}:\bm{\beta}]_{kl}.$
In SI Section V.2.2, we show that there exists $\bm{\beta}$, $\bm{\gamma}$ such that the Hermitian representation model (16) matches the ground truth Hamiltonian (in the CO basis).
II.2.3 Eight-Fold Symmetry-Preserving Regression
The Hermitian symmetry preserved by the first two models (11) and (16) can be viewed as a consequence of a deeper set of symmetries. Assuming real atomic orbitals $\{\phi_{j}\}$, Brailsford and Hall (1971) show that the $4$-index tensor $(ij|kl)$ defined in (5) satisfies the $8$-fold symmetry
Based on these observations, let us attempt to model the $4$-index tensor $(ij|kl)$ itself. The first part of our procedure is analogous to how we formed bases for the spaces of real symmetric and antisymmetric matrices, respectively, in Section II.2.2.
@@ II.2.3 Eight-Fold Symmetry-Preserving Regression
Here we form a basis $\mathscr{S}$ for the space of $N\times N\times N\times N$ tensors that satisfy the $8$-fold symmetry (17)—as shown in Brailsford and Hall (1971), this space has dimension
@@
To populate $\mathscr{S}$, we follow Algorithm 1. For $m=1,2,\ldots,8$, let $\sigma^{m}(i,j,k,l)$ be the $m$-th permutation seen in (17). For instance, $\sigma^{4}(i,j,k,l)=(l,k,j,i)$. Let $\mathcal{O}(i,j,k,l)$ denote $\{\sigma^{m}(i,j,k,l)\mid m=1,\ldots,8\}$, i.e., the orbit of the $4$-tuple $(i,j,k,l)$ under the group $\{\sigma^{m}\}_{m=1}^{8}$. In Algorithm 1, each value of $m$ corresponds to one unique orbit $\mathcal{O}(i,j,k,l)$. After Algorithm 1 terminates, $\mathscr{S}_{:,:,:,:,m}$ (letting the first four indices of $\mathscr{S}$ be free but fixing the last index at $m$) will be the $m$-th element of the basis, an $N\times N\times N\times N$ tensor that satisfies the symmetry (17). In the algorithm, the purpose of the binary tensor $\mathscr{B}$ is to signal whether we have already captured the orbit $\mathcal{O}(i,j,k,l)$. Upon the algorithm’s termination, $\mathscr{B}$ will be identically equal to $1$ and $m$ will equal $n_{T}$. With this framework, we build the model
@@
As $\tau$ is a linear combination of basis elements from $\mathscr{S}$, it is clear that $\tau$ will satisfy the $8$-fold symmetry (17). In this model, as long as $P$ is Hermitian, then $H$ will be Hermitian too:
@@
For the model (19), the parameters consist of $\bm{\beta}\in\mathbb{R}^{n_{T}}$. This is roughly $1/8$-th as many parameters as needed in the models (11) and (16).
In SI Section V.2.3, we explain how to compute $\bm{\beta}$ such that the model (19) matches the ground truth Hamiltonian.

### v2

@@ I Introduction
$\imath\partial_{t}\Psi(\mathbf{r},t)=\widehat{H}(\mathbf{r},t)\Psi(\mathbf{r},t).$ (1)
@@
In principle, time-dependent density functional theory (TDDFT) enables one to compute the reduced electron density $\rho$, without first solving the TDSE for $\Psi$ Ullrich (2012). However, in TDDFT, the density-dependent exchange-correlation potential (itself a part of the overall inter-electronic potential) is not known and must be approximated Gross and Maitra (2012); Maitra (2016). TDDFT is widely used to simulate electron dynamics and excited state properties, but in some cases the approximations employed for the exchange-correlation potential lead to inaccurate electron dynamics and charge transfer Elliott et al. (2012); Rappoport and Hutter (2012); Habenicht et al. (2014); Provorse, Habenicht, and Isborn (2015); Lacombe and Maitra (2020); Ranka and Isborn (2023). Learning a more accurate Hamiltonian, including the unknown density-dependent electron potential of TDDFT, would lead to advances in the simulation of the molecular properties and excited state dynamics.
@@
At the time of writing, there are only three works we know of that seek to learn the TDDFT exchange-correlation potential from data consisting of temporal snapshots of the electron density Suzuki, Nagai, and Haruyama (2020); Bhat et al. (2022); Yang and Whitfield (2023). These works focus on one-dimensional model problems; extending the methods to three-dimensional systems of chemical/physical interest will be a large-scale undertaking from both a theoretical and computational point of view. We lack an understanding of questions such as: (i) how should we form our training sets? (ii) how should we construct our model potentials? and (iii) how should we quantify errors? The literature on learning potentials for quantum molecular systems is tilted heavily toward static, time-independent problems Snyder et al. (2012); Bartók et al. (2017); Chandrasekaran et al. (2019); Schleder et al. (2019); Jørgensen et al. (2019); Smith et al. (2019); Ceriotti (2019) and offers little guidance on the answers to these questions, especially for molecular systems with, e.g., dozens of electrons.
@@
The broad goal of this paper is to extend our understanding of how to learn potentials from time series observations of quantum systems. We develop a learning framework for dynamics governed by the time-dependent Hartree-Fock (TDHF) equation. We focus on TDHF for four reasons. First, for TDHF, we know a priori how to compute the entire Hamiltonian, including the potential term we seek to learn. Hence we can compare our learned potentials against the ground truth. Second, TDHF is numerically tractable even for large molecular systems, so there is no a priori obstacle that stops us from generating training data. Third, Hartree-Fock theory incorporates an exact treatment of electron exchange. Consequently, the inter-electronic potential is purely a function of a $2$-electron integral tensor $(ij|kl)$—see (5)—that enjoys an $8$-fold permutation symmetry. Though existing density functionals that approximate electron exchange based on the local density do not have this symmetry, we expect an exact exchange-correlation functional to have symmetries similar to what we see in Hartree-Fock theory. This motivates studying how to incorporate symmetries into a machine learning model of a potential. Fourth, Hartree-Fock is itself the first model (approximating $\Psi$ using a single Slater determinant) in a sequence of increasingly accurate models that culminates in full configuration interaction (in which $\Psi$ is expanded using all possible Slater determinants excited from the Hartree-Fock reference determinant). Time-dependent full configuration interaction yields numerical solutions of (1) that are exact within a finite-dimensional basis set McWeeny (1989).
@@
Therefore, by studying and solving a machine learning problem for TDHF, we aim to produce insights that can be abstracted and generalized to learning problems that involve more accurate models for quantum dynamics. The present study aims to guide future efforts to learn a priori unknown terms that play important roles in time-dependent equations of motion. This includes, but is not limited to, learning the exchange-correlation potential for TDDFTSuzuki, Nagai, and Haruyama (2020); Bhat et al. (2022); Yang and Whitfield (2023), closure terms for time-dependent reduced density matrix functional theory and similar theoriesBlum (2012); Akbari et al. (2012); Ferré, Filatov, and Huix-Rotllant (2016); Bhat et al. (2024), and surrogate models for the Kadanoff-Baym equationsBassi et al. (2023).
@@
can be derived from the TDSE by approximating $\Psi$ as a single Slater determinant computed from a finite set of spin orbitals Szabo and Ostlund (2012). Here $P$ is the $1$-electron reduced density matrix corresponding to $\Psi$, while the Hamiltonian matrix $H$ is the operator $\widehat{H}$ from (1) expressed in a finite-dimensional basis McLachlan and Ball (1964); McWeeny (1989). Note that (2) only holds if one expresses $P$ and $H$ in orthonormal bases. The full, time-dependent Hamiltonian in (2) is
@@
On the right-hand side, the field-free Hamiltonian $H(P(t))$ includes the kinetic energy of the electrons, the potential due to interaction with the fixed nuclei, and the density dependent inter-electronic repulsion term, also known as the Fock matrix in Hartree-Fock theory. In (3), $V_{\text{ext}}(t)$ is an optional, external, time-dependent potential, e.g., an applied electric field. In our prior work Bhat, Ranka, and Isborn (2020); Gupta et al. (2022), we initiated the project of learning a model $\widetilde{H}(P)$ of the true field-free Hamiltonian $H(P)$ from time series observations $\{P(t_{j})\}$ of electron density matrices.
@@
Here we retain one feature of our prior work: we train our model $\widetilde{H}(P)$ on field-free trajectories, but we then test our models by quantifying how well they predict field-on dynamics. This is one way of measuring how well our models generalize to settings outside the training data. Further details are given in Sections II.3 and III.1.1.
In the present work, we move beyond prior work in three ways: (i) scaling our methods to much larger molecular systems while maintaining accuracy, (ii) training using high-quality ensembles (instead of single trajectories), and (iii) incorporating an $8$-fold symmetry that was ignored in prior work. We now explain these distinctions in turn.
First, we pay close attention to scalability to systems such as paranitroaniline ($\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$) in the STO-3G basis set, where the density matrices $P(t)$ are of size $60\times 60$. For $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$, the methods developed in this paper enable learning an inter-electronic potential parameterized by a tensor with more than $1.6$ million entries, with which we can propagate electron dynamics with an overall error on the order of $10^{-7}$—see Figure 3. This is six orders of magnitude better than what we can achieve using methods from our prior work. We are unaware of other methods that can scale to this system size. In the present work, we propose three models and derive expressions for Jacobian-vector products that enable matrix-free, iterative training. We formulate our models in such a way that they preserve symmetries using linear algebra. The combination of these techniques leads to scalability.
Second, in the present work, we train models using ensembles of field-free trajectories. In our past work, we only explored training on single field-free trajectories. We also pay close attention to the accuracy of these trajectories, employing a higher-order numerical method that, for the largest molecular systems we study, avoids spurious oscillations. We hypothesize that the ensemble more thoroughly explores the inter-electronic potential landscape. As we show, models trained on this ensemble have superior predictive power as measured by the ability to generate field-on trajectories $\widetilde{P}(t)$ that match true field-on trajectories $P(t)$.
Third, developing symmetry-constrained models lies at the core of the present work. For a variety of machine learning problems in the sciences, models that incorporate symmetries have achieved superior resultsZhang et al. (2018); Anderson, Hy, and Kondor (2019); Eismann et al. (2021); Satorras, Hoogeboom, and Welling (2021); Klus et al. (2021); Villar et al. (2021); Li et al. (2024); much of this literature focuses on continuous symmetry groups in real space. We present approaches tailored to the Hartree-Fock setting, but our methods can be generalized to other matrix/tensor-valued regression problems with permutation symmetries. Two of the models we propose preserve Hermitian symmetry. Another preserves a deeper $8$-fold permutation symmetry that has not been exploited in prior work. We express our model in a basis for the space of all $4$-index tensors that satisfy this symmetry, ensuring symmetry preservation.
@@ II.1 Choice of initial condition and time-stepping method
In order to train and test our models, we must have on hand a reliable set of electron dynamics trajectories. For a particular molecule and basis set, we apply standard electronic structure methods to compute the ground truth field-free Hamiltonian $H(P)$ and an initial condition $P(0)$ with an applied delta-kick perturbation. Before doing anything further, we evolve the density matrix forward in time to $P(2\delta t)$ with time step $\delta t=8.268\times 10^{-2}$ a.u. (atomic units). We then relabel this density matrix $P(2\delta t)$ as $P(0)$ and use this as our initial condition for all subsequent modeling/testing. We use a modified version of the Gaussian electronic structure code Frisch, Trucks, and et. al (2018) to compute the Hamiltonian $H(P)$ and the matrices $P(0)$ through $P(2\delta t)$.
To understand why we relabel $P(2\delta t)$ as $P(0)$, note that in our past Bhat, Ranka, and Isborn (2020); Gupta et al. (2022) and present work, our learning procedures depend on first estimating $dP/dt$ from $P(t)$ and then fitting $\imath dP/dt$ to $[H,P]$. That is, we minimize the residual between the left- and right-hand sides of (2). As expected given the nature of the Dirac delta, our numerical estimates of $dP/dt$ are much more accurate if we move past the time of the initial delta-kick perturbation. This improves all learning procedures we have tested.
Equipped with $H(P)$ and $P(0)$, our task is to numerically integrate (2) to produce a trajectory $\{P(t_{j})\}$ of electron density matrices on a temporal grid $t_{j}=j\Delta t$ for $j=0,\ldots,J$. In our prior work Bhat, Ranka, and Isborn (2020); Gupta et al. (2022), we employed the second-order modified midpoint unitary transformation (MMUT) method Li et al. (2005). In the present work, for reasons we now explain, we employ a fourth-order schemeCasas and Iserles (2006) that we abbreviate as CI4. We detail both schemes in Supplementary Material (SM) Section VI.1. Note that the CI4 method is based on a fourth-order Magnus expansion. In a comparative study of time-stepping methods for the time-dependent Kohn-Sham equations, a fourth-order Magnus integrator emerged as the clear winnerGémez Pueyo et al. (2018).
@@
Therefore, for all subsequent calculations, we use CI4Casas and Iserles (2006) to solve (2). This includes generation of training and test data using either ground truth Hamiltonians $H(P)$ or learned Hamiltonians $\widetilde{H}(P)$. We continue this thread in Section II.3 below.
@@ II.2 Linear models
We use $\widetilde{H}_{1}(P)$ to denote a model of the density-dependent, inter-electronic potential in (9). We present three scalable $\widetilde{H}_{1}(P)$ model, in increasing order of mathematical complexity. All three are linear in the entries of $P$. All three preserve Hermitian symmetry. The third model preserves an additional $8$-fold permutation symmetry that we detail below.
@@ II.2.1 Tied Regression
instead of (11b); for this model, the task is to learn both $\bm{\beta}$ and $\bm{\gamma}$. Model (11) effectively ties the parameters $\bm{\beta}$ and $\bm{\gamma}$ to one another (via $\bm{\beta}=\bm{\gamma}$), reducing the number of real parameters from $2N^{4}$ to $N^{4}$, and thus reducing training times. In SM Section VI.2.1, we show that there exists $\bm{\beta}$ such that the tied model (11) matches the ground truth Hamiltonian (in the CO basis). Thus there is nothing gained by considering a model in which (12) replaces (11b).
@@ II.2.2 Regression with Hermitian Representation
$\sum_{i,j}P^{R}_{ij}\beta_{ijkl}=[P^{R}:\bm{\beta}]_{kl}.$
In SM Section VI.2.2, we show that there exists $\bm{\beta}$, $\bm{\gamma}$ such that the Hermitian representation model (16) matches the ground truth Hamiltonian (in the CO basis).
II.2.3 Eight-Fold Symmetry-Preserving Regression
The Hermitian symmetry preserved by the first two models (11) and (16) is a consequence of a deeper set of symmetries. Assuming real atomic orbitals $\{\phi_{j}\}$, the $4$-index tensor $(ij|kl)$ defined in (5) satisfiesBrailsford and Hall (1971) the $8$-fold symmetry
Based on these observations, let us attempt to model the $4$-index tensor $(ij|kl)$ itself. The first part of our procedure is analogous to how we formed bases for the spaces of real symmetric and antisymmetric matrices in Section II.2.2.
Here we form a basis $\mathscr{S}$ for the space of $N\times N\times N\times N$ tensors that satisfy the $8$-fold symmetry (17)—this space has dimensionBrailsford and Hall (1971)
@@ II.2.3 Eight-Fold Symmetry-Preserving Regression
To populate $\mathscr{S}$, we follow Algorithm 1. For $m=1,2,\ldots,8$, let $\sigma^{m}(i,j,k,l)$ be the $m$-th permutation seen in (17). For instance, $\sigma^{4}(i,j,k,l)=(l,k,j,i)$. Let $\mathcal{O}(i,j,k,l)$ denote $\{\sigma^{m}(i,j,k,l)\mid m=1,\ldots,8\}$, i.e., the orbit of the $4$-tuple $(i,j,k,l)$ under the group $\{\sigma^{m}\}_{m=1}^{8}$. In Algorithm 1, each value of $m$ corresponds to one unique orbit $\mathcal{O}(i,j,k,l)$. After Algorithm 1 terminates, $\mathscr{S}_{:,:,:,:,m}$ (letting the first four indices of $\mathscr{S}$ be free but fixing the last index at $m$) will be the $m$-th element of the basis, an $N\times N\times N\times N$ tensor that satisfies the symmetry (17). In the algorithm, the purpose of the binary tensor $\mathscr{B}$ is to signal whether we have already captured the orbit $\mathcal{O}(i,j,k,l)$. Upon the algorithm’s termination, $\mathscr{B}$ will be identically equal to $1$ and $m$ will equal $n_{T}$. With this framework, we build the $8$-fold symmetry-preserving model
@@
As $\tau$ is a linear combination of basis elements from $\mathscr{S}$, it is clear that $\tau$ will satisfy the $8$-fold symmetry (17). In this model, as long as $P$ is Hermitian, so is $H$:
@@
For the model (19), the parameters consist of $\bm{\beta}\in\mathbb{R}^{n_{T}}$. This is roughly $1/8$-th as many parameters as needed in the models (11) and (16). In SM Section VI.2.3, we derive $\bm{\beta}$ such that the model (19) matches the ground truth Hamiltonian.